# Towards Aligning Multimodal LLMs with Human Experts: A Focus on Parent–Child Interaction

### v1

@@ 3.6.2. Illustrative Examples of Diverging SLP Judgements
SLP1 adopted a balanced and multimodal approach. While gaze was still considered, it was treated as one of several contributing cues–alongside gestures, verbal referencing, emotional affect, and body orientation. SLP2 often acknowledged child-led play and nonverbal forms of engagement, showing flexibility in how joint attention could be expressed. Segments with inconsistent gaze but strong affective presence were still given high ratings when the child’s overall behaviour reflected social coordination.
SLP2, in contrast, placed strong emphasis on visual engagement, particularly gaze alternation between the adult and shared object. Eye contact and sustained visual referencing were seen as essential for identifying strong joint attention. In the absence of clear gaze cues, even segments involving verbal speech or physical interaction were often rated as moderate or poor. For example, SLP1 rated several segments as poor when the child appeared engaged but did not meet the baseline requirement of mutual gaze.
@@
Together, these examples demonstrate how gaze, vocalisation, and action were weighted differently by each expert, and how their background and theoretical orientation shaped their interpretation
@@ 4.1.1. Method
With the rapid advancement of MLLMs, their application to human behaviour analysis presents a promising frontier. In this study, we adopt Gemini 2.5 Pro [^5] https://blog.google/products/gemini/gemini-2-5-pro-updates/ for its strong multimodal reasoning capabilities (yue2024mmmu; team2024gemini; fu2024video). However, prior research (shi2025towards) shows that directly prompting MLLMs to identify key interaction segments–especially for complex, socially embedded tasks like joint attention–often fails. Models continue to struggle with long video understanding, temporal grounding, and fine-grained cues such as gaze interpretation.
@@
Structured Behaviour Description Prompt (Gaze-Action-Vocalisation):
@@ 4.1.2. Result
To evaluate the MLLM’s behavioural description performance, we compared its outputs with reference annotations created by our research team. These reference annotations were constructed in two stages, following the same gaze-action-vocalisation structure used in the prompt. Importantly, our annotation process was informed by the interview findings with SLPs–particularly their attention to detail and tendency to distinguish subtle variations in gaze direction, gesture context, and vocalisation specificity.
For each segment, we began by manually reviewing the MLLM-generated outputs. In the first stage, we corrected instances where the generated description contradicted the observed behaviour–for example, changing “the child looked at the parent” to “the child looked down at the table” if no eye contact was actually made. In the second stage, we refined underspecified descriptions by aligning them with SLP-like granularity–for instance, revising “the child looked at the parent” to “the child looked at the parent’s hand” when the visual target was more specific. These two-step corrections ensured that reference labels reflected both factual accuracy and expert-informed attentional focus.
Based on this corrected reference set, we computed segment-level accuracy scores for each of the three behavioural fields. As shown in Table 4, the MLLM performed best in the action category, with a mean accuracy of 0.88, followed by vocalisation (0.87) and gaze (0.86). All three fields achieved perfect accuracy in at least one video, but gaze also had the lowest minimum, highlighting its difficulty for the model.
Table 4. Accuracy statistics by behavioural field
@@
These results demonstrate that our structured prompting approach substantially improved the reliability of MLLM-generated behavioural annotations, enabling the model to produce interpretable outputs that aligned with expert-labelled segment judgments in many cases. Compared to Shi et al. (shi2025towards), where average behavioural and gaze accuracy was around 62%, our structured prompting approach led to substantial improvements in both description accuracy and eye-contact interpretation.
@@ 4.2.1. Dataset
We assembled a text-based agreement dataset from 25 parent–child interaction (PCI) videos analysed in Stage 1. In this stage, each segment was coded along three behavioural cue dimensions–gaze, action, and vocalisation–for both the parent and the child using AI. We reviewed the Stage 1 AI annotations of behaviour, corrected erroneous information, and consolidated them into $(\textit{parent, child}),\textit{label}$ pairs. Only segments with at least two SLPs in agreement were retained, and all cases of complete disagreement were removed. This resulted in a total of 615 agreed pairs.
@@ 4.2.2. Experiments
We employed GPT-4.1 (gpt-4.1-2025-04-14) [^6] https://platform.openai.com/docs/models/gpt-4.1–at present OpenAI’s most capable non-reasoning model–as the backbone model for text-based judgements. Evaluation was conducted on a video-by-video basis, where all segments from the same video were included together in the prompt to preserve contextual continuity. Each segment was represented as a structured observation in the form of paired $(\textit{parent},\textit{child})$ descriptions across gaze, action, and vocalisation, and the model was required to assign one overall label: Strong, Moderate, or Poor.
@@
In the many-shot condition, we adopted a leave-one-video-out strategy (zhao2024learning): the held-out video served as the evaluation set, while annotated examples from all remaining videos-each consisting of $(\textit{parent},\textit{child})$ observations paired with a gold label-were embedded in the prompt as demonstrations. This design ensured that the model had access to the full range of SLP-provided knowledge from other videos when making judgements on the target segments. The procedure was repeated across all videos, so that every video was evaluated once as the hold-out, and the union of predictions from all folds yielded the complete set of model-generated judgement labels.
@@ 4.2.4. Result: Many-shot Prompting Strategy Yields Better Alignment
Many-shot prompting led to stronger alignment with expert consensus than zero-shot prompting (see Figure 5). Accuracy rose from 0.44 to 0.57, and Cohen’s $\kappa$ nearly doubled (0.10 → 0.18), indicating more consistent agreement beyond chance. Macro-precision also improved slightly (0.40 → 0.42), suggesting fewer false positives, while macro-recall decreased (0.53 → 0.46), implying a more conservative stance that missed some expert-labelled cases. Despite this trade-off, the overall macro-F1 increased from 0.34 to 0.40, reflecting a net gain in balanced performance.
Per-class performance: Strong and Moderate improve, Poor remains unreliable.
At the category level, many-shot prompting yields improvements for Strong and Moderate cases but not for Poor. For Strong, recall rises (0.48→0.60) and F1 improves (0.41→0.49), indicating more reliable detection of clear joint attention. For Moderate, which dominates the dataset, both precision (0.80→0.83) and recall (0.42→0.57) increase, producing the largest F1 gain (0.55→0.67). By contrast, Poor segments remain highly challenging: recall drops (0.70→0.20) and F1 declines (0.07→0.04). This reflects two issues. First, zero-shot prompting severely over-predicted poor cases, inflating recall but with almost no precision (0.04). many-shot prompting suppresses this over-prediction, yielding more conservative outputs but still very low precision (0.02). Second, the data distribution itself is highly imbalanced: across the dataset, Moderate accounts for 469 segments (73.5%), Strong for 136 (21.3%), while Poor is represented by only 10 examples (1.6%). Such scarcity makes it difficult to validate or learn reliable patterns for low-fidelity behaviours in a many-shot condition.
Figure 5. Radar plots comparing zero-shot (red) and many-shot (blue) prompting across overall and per-class evaluation metrics. many-shot consistently improves accuracy, macro-F1, and Cohen’s $\kappa$ in the overall condition, while also boosting performance on Strong and Moderate categories. Performance on the Poor category remains weak under both conditions, reflecting its limited representation in the dataset.:
@@ 5.1. Greater Alignment at the Level of Behavioural Observations
SLPs showed strong alignment with one another with behavioural observation, consistently pointing to gaze, action, and vocalisation as the central dimensions for describing joint attention. Our structured prompting method leveraged this shared consensus and achieved similarly high alignment: MLLMs could reliably capture and reproduce these behavioural cues.
Because expert criteria were already aligned, observation proved to be a tractable entry point for human–AI alignment. By guiding MLLMs to “see” behaviours through these shared dimensions, we can achieve stable and reproducible outputs. This suggests that systems for analysing Parent-Child Interaction should prioritise clear, structured observation before attempting higher-level reasoning.
The relative ease of alignment in observation highlights a broader design lesson: when experts already share common descriptors, MLLMs can be scaffolded to replicate them reliably. Multimodal models are particularly well-suited to capture low-level, directly perceivable cues—for instance, gaze direction, vocal onset, or gesture trajectories—because such behaviours are already described consistently across experts. Achieving alignment in these cases mainly requires iterative confirmation with human experts, rather than inventing new criteria.
For example, in education (xing2024survey), teachers can consistently agree on whether a child raised a hand, spoke out loud, or looked at the teacher. In UX evaluation (vermeeren2010user), experts reliably agree on whether a user paused, clicked, or hovered over an interface element. In creative work critiques (alabood2023systematic), reviewers may all notice surface-level features such as colour saturation or the use of bold lines. These cues are easier for multimodal models to observe, making observation alignment relatively tractable.

### v2

@@ 3.6. Annotation Study Results: SLPs’ Observational Judgements of Joint Attention
To quantify the level of consistency among the three experts, we computed inter-annotator agreement using Krippendorff’s $\alpha$ with an ordinal scale, which reflects the ordered nature of the Strong, Moderate, and Poor labels. The overall $\alpha$ was 0.19, indicating low consistency among the three SLPs. This value indicates low consistency among the three SLPs. In combination with the heavily skewed label distribution, where each of the three experts marked more than 60% of segments as Moderate, the result suggests that disagreements often arose in borderline cases rather than in clearly distinct behaviours.
Pairwise analyses further reveal asymmetries in expert alignment. The highest agreement occurred between SLP1 and SLP2 ($\alpha=0.31$), while substantially lower agreement was observed for SLP1–SLP3 ($\alpha=0.15$) and SLP2–SLP3 ($\alpha=0.11$). These differences suggest that the three SLPs do not form a single coherent “consensus group”; instead, their judgments cluster into at least two partially aligned perspectives. Taken together, the low overall $\alpha$ and uneven pairwise consistency indicate that expert interpretation of joint attention is inherently variable, especially in segments where behaviours fall between categories. This variability reinforces the importance of modelling expert diversity rather than assuming a single ground truth.
@@ 3.6.2. Illustrative Examples of Diverging SLP Judgements
SLP1 adopted a balanced and multimodal approach. While gaze was still considered, it was treated as one of several contributing cues–alongside gestures, verbal referencing, emotional affect, and body orientation. SLP1 often acknowledged child-led play and nonverbal forms of engagement, showing flexibility in how joint attention could be expressed. Segments with inconsistent gaze but strong affective presence were still given high ratings when the child’s overall behaviour reflected social coordination.
SLP2, in contrast, placed strong emphasis on visual engagement, particularly gaze alternation between the adult and shared object. Eye contact and sustained visual referencing were seen as essential for identifying strong joint attention. In the absence of clear gaze cues, even segments involving verbal speech or physical interaction were often rated as moderate or poor. For example, SLP2 rated several segments as poor when the child appeared engaged but did not meet the baseline requirement of mutual gaze.
@@
Together, these examples demonstrate how gaze, vocalisation, and action were weighted differently by each expert, and how their background and theoretical orientation shaped their interpretation. This variability in interpretive emphasis directly aligns with the earlier statistical results, where overall inter-annotator agreement remained low ($\alpha=0.19$), reflecting the inherent subjectivity of these judgements.
@@ 4.1.1. Method
With the rapid advancement of MLLMs, their application to human behaviour analysis presents a promising frontier. In this study, we adopt Gemini 2.5 Pro [^5] https://blog.google/products/gemini/gemini-2-5-pro-updates/ for its strong multimodal reasoning capabilities (Yue et al., 2024; Team et al., 2024; Fu et al., 2025). A further practical consideration is that Gemini 2.5 Pro is, at present, the only widely accessible API model that accepts direct video inputs. Other foundation models, including the latest OpenAI offerings, currently support only image-based visual inputs, which makes them less suitable for segment-level behavioural captioning in continuous parent-child interaction videos. However, prior research (Shi et al., 2025) shows that directly prompting MLLMs to identify key interaction segments–especially for complex, socially embedded tasks like joint attention–often fails. Models continue to struggle with long video understanding, temporal grounding, and fine-grained cues such as gaze interpretation.
@@
Structured Behaviour Description Prompt
(Gaze-Action-Vocalisation)
@@ 4.1.2. Result
To evaluate the MLLM’s behavioural description performance, we first conducted a structured human review of all model-generated outputs for each segment. Two annotators from our research team independently examined every 5-second segment. For each segment, they reviewed all six model-generated entries (parent/child $\times$ gaze/action/vocalisation) and assigned a binary correctness label: 1 if the description accurately matched the observed behaviour in the video, and 0 otherwise. For each of the three behavioural fields, correctness was determined according to the subject–verb–object structure specified in the prompt. For gaze, an entry was marked as correct only if both the subject and the gaze target (object) were accurate, such as distinguishing between looking at the parent’s face or looking at the blocks held in the parent’s hand. For action, correctness required that the subject and the action described were consistent with the video, for example identifying whether the child was using a hand to pick up a block or performing another distinct motor behaviour. For vocalisation, an entry was considered correct only when the described verbal content matched what was actually spoken in the segment, including both who produced the speech and the approximate content of the utterance. These field-specific criteria ensured that the binary labels reflected the accuracy of the subject–verb–object components rather than surface-level wording.
The two annotators achieved an inter-annotator agreement of Cohen’s $\kappa=0.92$, indicating high consistency. Any disagreements were resolved through discussion, and most arose from fine-grained distinctions in gaze descriptions. For example, the model sometimes output “child look at parent” when the behaviour should instead specify whether the child was looking at the parent’s face or hand. Based on these human-verified binary correctness labels, we computed accuracy for each behavioural field at the level of entire videos. For each video, we aggregated all segment-level binary labels for that field and calculated a per-video accuracy score. Table 4 reports the mean, median, maximum, and minimum of these per-video accuracies across the full dataset. As shown in the table, the MLLM performed best in the action category, with a mean per-video accuracy of 0.88, followed by vocalisation (0.87) and gaze (0.86). All three fields achieved perfect accuracy in at least one video, although gaze also had the lowest minimum accuracy, highlighting its difficulty for the model.
Table 4. Per-video aggregated accuracy statistics. For each field, accuracies were computed per video, and the table summarises the mean, median, maximum, and minimum across all videos.
@@
These results demonstrate that our structured prompting approach substantially improved the reliability of MLLM-generated behavioural annotations, enabling the model to produce interpretable outputs that aligned with expert-labelled segment judgments in many cases. Shi et al. (Shi et al., 2025) first introduced this behavioural description task and dataset, reporting that the model achieved approximately 62% behavioural accuracy within the 23% of segments where it successfully detected gaze cues. Building on this prior setup, our structured prompting approach ensures that the model receives explicit action, vocalisation, and gaze information for every segment and thus produces complete descriptions for 100% of the dataset. Using a comparable combined accuracy metric that marks a segment as correct only when all three dimensions are correct, our method achieved 81% accuracy. This reflects a substantial improvement in both overall description quality and gaze-related interpretation.
@@ 4.2.1. Dataset
The Stage 2 dataset was derived from the model-generated segment descriptions produced in Stage 1. For each 5-second segment across the 25 parent–child interaction videos, the model output six behavioural entries (parent/child $\times$
gaze/action/vocalisation). Two annotators independently evaluated the correctness of each entry based on the video, and any disagreements were resolved through discussion until consensus was reached. This consensus served as the ground truth for determining which descriptions required correction.
We then applied targeted revisions only to the entries marked as incorrect during the agreement process. Following a minimal-editing principle, we retained the original subject–verb–object structure whenever possible and corrected only the behavioural element that was factually inaccurate or mis-specified (e.g., replacing “the child looked at the parent” with “the child looked down at the table” if no eye contact occurred). When the model’s description was underspecified but directionally correct, we refined it to align with SLP-like granularity, such as specifying the actual attentional target (e.g., “the child looked at the parent’s hand”).
The resulting set of minimally corrected and consensus-verified descriptions constituted the input to the Stage 2 judging process. [^6] We will release the anonymised annotated dataset and analysis materials upon acceptance to support transparency and reproducibility.
@@ 4.2.2. Experiments
For the Stage 2 judging process, we used GPT-4.1 (gpt-4.1-2025-04-14) [^7] https://platform.openai.com/docs/models/gpt-4.1 as the backbone model for all text-based decisions. Stage 2 mirrors how SLPs conduct holistic assessment: instead of observing isolated 5-second segments, the model must read the entire sequence of segment-level descriptions from a full video and determine which portions correspond to Strong, Moderate, or Poor joint attention. This requires integrating behavioural evidence distributed across the interaction, similar to how an SLP reviews the full temporal flow of gaze, action, and vocalisation before forming a judgement. The task therefore primarily requires stable long-context understanding, enabling the model to integrate behavioural descriptions that span the full interaction rather than treating segments independently.
Although no existing LLM-as-judge benchmark directly evaluates this form of segment-level behavioural scoring, the Stage 2 task imposes requirements that differ substantially from Stage 1. Stage 1 relies on Gemini 2.5 Pro, a strong video–reasoning model that excels at extracting behavioural cues from short, self-contained 5-second clips. In contrast, Stage 2 mirrors how SLPs make holistic judgements: the model must read the entire sequence of segment descriptions for a full video, integrate behavioural evidence distributed across many segments, and determine which portions correspond to Strong, Moderate, or Poor joint attention. In our dataset, the longest 12-minute video produced approximately 60,000 tokens of segment-level descriptions, meaning the Stage 2 model must process and integrate an input roughly of this size for a single judgement. As a result, long-context reliability, rather than step-by-step reasoning, becomes the primary requirement for Stage 2.
At the time of our experiments, GPT-4.1 was the strongest publicly available model for long-context understanding. It supports a verified 1M-token context window and demonstrates superior performance compared with GPT-4o and o3 models on long-context retrieval and multi-hop long-context reasoning [^8] https://openai.com/index/gpt-4-1/. These benchmarks indicate that GPT-4.1 is particularly reliable at retrieving, tracking, and integrating information scattered across extended contexts. Such characteristics align directly with the needs of Stage 2, where the model must synthesise behavioural descriptions across an entire video without speculative inference. For these reasons, we selected GPT-4.1 as the backbone for Stage 2 rather than reusing the Stage 1 model or employing a specialised reasoning model.
Evaluation in Stage 2 followed the same holistic procedure used by SLPs when annotating parent–child interactions. All segments from the same video were provided together in a single prompt, allowing the model to review the full sequence of behaviours before making any judgement. Each segment was represented as a structured pair of $(\textit{parent},\textit{child})$ observations across gaze, action, and vocalisation. Rather than assigning labels to segments independently, the model was required to integrate information across the entire video and identify which parts of the interaction corresponded to Strong, Moderate, or Poor joint attention, mirroring the way SLPs derive their final ratings from the full interaction.
@@
In the many-shot condition, we adopted a leave-one-video-out strategy (Zhao et al., 2024): the held-out video served as the evaluation set, while annotated examples from all remaining videos-each consisting of $(\textit{parent},\textit{child})$ observations paired with a gold label-were embedded in the prompt as demonstrations. This design ensured that the model had access to the full range of SLP-provided knowledge from other videos when making judgements on the target segments. The procedure was repeated across all videos, so that every video was evaluated once as the hold-out, and the union of predictions from all folds yielded the complete set of model-generated judgement labels.
@@
For both the zero-shot and many-shot settings, we evaluated the model using three decoding temperatures (0, 0.3, and 0.5). Because the judgment task requires stable and reliable decisions rather than highly diverse generations, we intentionally restricted the temperature range to relatively low values. To account for the stochasticity introduced at temperatures above zero, we conducted three independent runs for each non-deterministic setting (0.3 and 0.5). Temperature 0 is deterministic and therefore required only a single run.
@@ 4.2.4. Result: Many-shot Prompting Strategy Yields Better Alignment
Temperature effects.
Figure 5 shows clear temperature-dependent differences in model reliability across both prompting modes. Temperature 0.3 consistently produced the strongest and most stable performance: in the many-shot condition it achieved the highest accuracy ($0.64\pm 0.01$), macro-F1 ($0.44\pm 0.01$), and Cohen’s $\kappa$ ($0.26\pm 0.02$), while zero-shot prompting likewise reached its best scores at this setting (macro-F1 $0.351\pm 0.005$, $\kappa=0.103\pm 0.006$). In contrast, temperature 0 underperformed in both modes (e.g., many-shot macro-F1 =0.40=0.40, zero-shot macro-F1 =0.344=0.344). Temperature 0.5 yielded slightly lower means and substantially higher variance; for example, in many-shot prompting accuracy dropped to $0.61\pm 0.05$ and macro-F1 to $0.42\pm 0.02$, suggesting that excessive sampling randomness reduces stability. Overall, temperature 0.3 offers the best balance of accuracy and robustness, and we therefore adopt it as the default for subsequent analyses.
Figure 5. Effect of decoding temperature on model reliability for zero-shot and many-shot prompting. Each subplot shows mean performance with standard-deviation error bars across temperatures 0, 0.3, and 0.5 for accuracy, macro-precision, macro-recall, macro-F1, and Cohen’s $\kappa$. Across all metrics, temperature 0.3 provides the most reliable overall performance, while temperature 0 underperforms and temperature 0.5 exhibits reduced stability due to increased sampling randomness. Many-shot prompting benefits more from moderate stochasticity than zero-shot prompting, which remains comparatively less sensitive to temperature changes. These results support the choice of temperature 0.3 for subsequent analyses.
Five side-by-side line plots compare accuracy, macro-precision, macro-recall, macro-F1, and Cohen’s kappa across temperatures 0, 0.3, and 0.5 for zero-shot and many-shot prompting. Temperature 0.3 yields the most stable and consistent performance, temperature 0 performs worse overall, and temperature 0.5 shows higher variance. Many-shot prompting is more affected by temperature changes than zero-shot prompting.
@@
Many-shot prompting achieved stronger alignment with expert consensus than zero-shot prompting (see Figure 6). Accuracy improved from 0.45 to 0.64, and Cohen’s $\kappa$ increased from 0.10 to 0.26. While Cohen’s $\kappa$ shows an improvement, it still remains within the slight-agreement range. Macro-precision also rose modestly (0.40 → 0.44), indicating fewer false positives, while macro-recall decreased slightly (0.54 → 0.50), suggesting a more conservative stance that misses some expert-labelled cases. Despite this trade-off, the overall macro-F1 increased from 0.35 to 0.44, representing a clear improvement in balanced performance.
Per-class performance.
At the category level, many-shot prompting yields improvements for Strong and Moderate cases but not for Poor. For Strong, recall increases from 0.47 to 0.64 and F1 improves from 0.41 to 0.52, indicating more reliable detection of clear joint attention. For Moderate, which dominates the dataset, both precision (0.80→0.85) and recall (0.44→0.65) increase, producing the largest F1 gain (0.57→0.74). By contrast, Poor segments remain highly challenging: recall drops sharply (0.70→0.20) and F1 decreases (0.07→0.06). This reflects two issues. First, zero-shot prompting substantially over-predicts Poor cases, producing high recall but extremely low precision (0.04). Many-shot prompting suppresses this over-prediction, yielding more conservative outputs but still very low precision (0.04→0.04). Second, the dataset distribution is highly imbalanced: Moderate accounts for 469 segments (73.5%), Strong for 136 (21.3%), and Poor for only 10 examples (1.6%). Such scarcity makes it difficult for the model to learn or validate reliable patterns for low-fidelity Poor behaviours, regardless of prompting method.
Figure 6. Radar plots comparing zero-shot (red) and many-shot (blue) prompting across overall and per-class evaluation metrics. many-shot consistently improves accuracy, macro-F1, and Cohen’s $\kappa$ in the overall condition, while also boosting performance on Strong and Moderate categories. Performance on the Poor category remains weak under both conditions, reflecting its limited representation in the dataset.
The radar plot compares the performance of zero-shot (red) and many-shot (blue) prompting, showing that many-shot prompting consistently improves overall metrics and performance for the ’Strong’ and ’Moderate’ categories, while performance on the ’Poor’ category remains low due to its limited dataset representation.
@@ 5.1. Greater Alignment at the Level of Behavioural Observations
SLPs showed strong alignment with one another in behavioural observation, consistently pointing to gaze, action, and vocalisation as the central dimensions for describing joint attention. Our structured prompting method leveraged this shared consensus and achieved similarly high alignment. MLLMs could reliably capture and reproduce these behavioural cues. Because experts already converge on these descriptors, observation proved to be a tractable entry point for human–AI alignment.
This tractability reflects a broader phenomenon. When the domain involves low-level, directly perceivable behaviours, experts tend to describe events in highly similar ways. In education (Xing et al., 2024), teachers reliably agree on whether a child raised a hand or spoke out loud. In UX evaluation (Vermeeren et al., 2010), experts consistently identify pauses, clicks, and hovers. In creative work critiques (Alabood et al., 2023), reviewers often converge on surface-level cues such as colour saturation or bold outlines. These shared descriptors create a stable grounding that multimodal models can be scaffolded to replicate. As a result, observation-level alignment can be achieved before attempting any higher-level judging.
This suggests that systems for analysing Parent-Child Interaction should prioritise clear and structured observation as their first layer of alignment. By guiding MLLMs to “see” behaviours through common expert descriptors, we can produce stable and interpretable outputs that form a reliable substrate for subsequent judging tasks.
Design Takeaway #1: Provide Perceptual Assistance With Transparent and Interpretable Cues. MLLMs can meaningfully support early-stage observation. In our study, the model achieved 81% accuracy in identifying low-level behavioural cues, showing strong alignment with SLPs in detecting gaze shifts, object-directed actions, and vocalisations. However, these machine-generated observations cannot replace human observation, especially in a clinically serious and nuanced behavioural assessment task. Even with relatively strong cue-level performance, SLPs should not simply trust model outputs. They must actively review, verify, and integrate these cues into their own observation workflow.
These cues nevertheless remain valuable. They are precisely the behaviours that SLPs spend substantial time documenting during and after sessions, both for clinical records and to support later judgement. SLP2 specifically noted that gathering concrete behavioural examples is a time-consuming but essential part of subsequent assessment. By surfacing structured and timestamped observations, MLLMs can reduce documentation burden and provide a consistent baseline for longitudinal tracking, provided that SLPs remain the primary interpreters.
This highlights a concrete design direction. Systems should help SLPs use model-extracted cues efficiently and reflectively. Interfaces should allow interactive editing of cue descriptions, timestamp correction, merging or splitting of events, and clinician-authored annotations. These features can transform model detections into editable professional artefacts. Transparent visualisations, such as layered timelines, collapsible cue categories, uncertainty indicators, and explicit evidence, can further support rapid verification while preventing over-reliance on automation. Through these mechanisms, perceptual assistance augments rather than replaces human expertise and embeds MLLM-derived cues into clinical judgement in a controlled and interpretable manner.